Please help.
SAT Problems:
For what values of p does the equation px = 10 have a solution −5; 1; 20?
For what values of m does the equation mx = 5 have a single solution? Can m have a value for which the equation does not have solutions; have infinite number of solutions?
Miriam left Chicago at 9:00 AM and headed to Washington. The first 2 hour Miriam averaged 70 mph. But after she decided to reach Washington at 6:00 PM, she increased her speed. What was her speed after increasing, if the distance between the cities is 700 miles.
I actually tried to solve these problems. Please help(please include the steps too). Thank you! Thank you so much!

Answer by greenestamps(13209)



For the first two questions, the statements of the problems are not at all clear....

For any equation of the form A*B=C, as long as C is not 0, the equation has a single solution for A given a value for B, and a single solution for B given a value of A.

If in an equation of that form C=0, then one of A or B has to be 0; then there is an infinite number of solutions for the other number, since any number times 0 is 0.

There are no equations of that form that have no solutions.

last question....

The trip is 9 hours, from 9am to 6pm; in the first 2 hours at 70mph, she travels 70*2=140 miles; she has 700-140=560 miles left to go; she wants to travel those 560 miles in 9-2=7 hours.

What speed is required to travel 560 miles in 7 hours?